# Quantum-enhanced Markov Chain Monte Carlo for systems larger than your Quantum Computer

### v1

@@ Appendix A Implementation details
The numerical quantum simulations were carried out using python 3.11, running qulacs quantum circuit simulator [31] and can be found open source on Github [32]. The base of the code was built upon a pre-existing library, however it was drastically modified to fit the requirements of this paper [33].
To keep consistent with the original work, parameters have been kept constant as much as possible [1]. For example, $\gamma$ is varied in the range $(0.25,0.6)$ and the integer time parameter is varied between $(2,20)$ for all of the following experiments. One key difference is the method of calculating the spectral gap for quantum proposal methods. The authors of [1] employed exact numerical integration by discretising the gamma parameter. Here however, we employ Monte-Carlo integration by uniformly sampling the $t$ and $\gamma$ regions. Explicitly, this means that for a given row of the $Q$ matrix, the measurement statistics are estimated by averaging over $30$ (a value chosen by experimentation) randomly generated $t$ and $\gamma$ values. When one is doing plain QeMCMC or CGQeMCMC without multiple groups, the $i^{th}$ row of $Q$ can be found by doing the following 30 times: First, randomly generate $t$ and $\gamma$ values then use as quantum input the $i^{ith}$ state, run the appropriate quantum circuit and output the statevector. Clearly, averaging over the 30 statevectors gives a good approximation of the likely transitions between states.
@@ B.1 Hamming distance and Energy difference of proposals
In this section, the Hamming distance and energy difference of proposed steps are compared. The classical local step will only ever take steps of Hamming distance $=1$, meaning it is entirely local but has a high acceptance rate as the energy difference of proposed steps is low. This means that such updates traverse the cost function landscape in a very local way, often getting stuck in local minima for very long periods of time. The uniform proposal has the extreme opposite feature, as it picks a possible step at random, it has a very low acceptance rate meaning slow exploration of the landscape as the energy of a proposed state is unlikely to be close or lower. When a step is accepted however, the hamming distance is very large. Both of the above points are made clear in Figure 6.
Any CG method that does not multiple sample will have the locality issue of occasionally getting stuck in minima. If one does multiple sample, it becomes clear that the Hamming distance is slightly reduced compared to the uniform or QeMCMC method. The energy difference of proposals still outperforms classical methods, almost matching the QeMCMC. One interesting artefact is that the improved local group (without multiple sampling) greatly reduces the energy difference compared to all methods, while increasing the Hamming distance compared to the classical local method. This means that if one were to choose a problem where the local method performs well (for example in particular temperature regimes) then the improved local method is always likely to sustainably outperform the classical proposals. Scaling of this artefact is left to future work.
@@ B.2 Temperature and convergence for large systems
The relationship between temperature, system size and thermalisation is conveyed through analysis of the energy convergence in Figure 7. One can see that the thermalisation of the uniform proposal becomes useless for large systems, the local proposal struggles at low temperatures and in every case, both quantum proposals analysed efficiently find a better than classical low energy regime. Magnetisation plots are not provided as they falsely suggest convergence of non-converged distributions.
Figure 7: Average energy thermalisation plot of 10 Markov Chains for one randomly selected Ising model of size $16$ (left), $25$ (middle) and $36$ (right). $T=0.1$ (top) is clearly a more difficult problem than $T=1$ (bottom) where the Markov chains acceptance rate is much higher. The coloured dotted lines represent the $10$ lowest energy levels of the system, while the black line represents the average Boltzmann energy. This is not provided for the 36 spin system due to the computation limitations of brute force solvers.
@@ B.3 Temperature and scaling
The temperature dependence of spectral gap scaling is briefly conveyed here through a cross-section of temperature regimes. Figure 8 shows the scaling behavior of each proposal method for $T=\{0.1,1,10\}$ while Table 1 shows results of each fit according to Equation 15.

### v2

@@ Appendix A Implementation details
The numerical quantum simulations were carried out using python 3.11, running qulacs quantum circuit simulator Suzuki et al. (2021) and can be found open source on Github Ferguson (2024). The base of the code was built upon a pre-existing library, however it was drastically modified to fit the requirements of this paper Pal (2023).
To keep consistent with the original work, parameters have been kept constant as much as possible Layden et al. (2023). For example, $\gamma$ is varied in the range $(0.25,0.6)$ and the integer time parameter is varied between $(2,20)$ for all of the following experiments. One key difference is the method of calculating the spectral gap for quantum proposal methods. The authors of Layden et al. (2023) employed exact numerical integration by discretizing the gamma parameter. Here however, we employ Monte Carlo integration by uniformly sampling the $t$ and $\gamma$ regions. Explicitly, this means that for a given row of the $Q$ matrix, the measurement statistics are estimated by averaging over $m$ (chosen to be 30 by experimentation) randomly generated $t$ and $\gamma$ values. When one is doing plain QeMCMC or CGQeMCMC without multiple groups, the $i^{th}$ row of $Q$ can be found by doing the following $m$ times: First, randomly generate $t$ and $\gamma$ values then use as quantum input the $i^{th}$ state, run the appropriate quantum circuit and output the statevector. Averaging over the $m$ statevectors gives a good approximation of the likely transitions between states.
@@ B.1 Hamming distance and Energy difference of proposals
In this section, the Hamming distance and energy difference of proposed steps are compared through Figure 6. The classical local step will only ever take steps of Hamming distance $=1$, meaning it is entirely local but has a high acceptance rate as the energy difference of proposed steps is low. This means that such updates traverse the cost function landscape in a very local way, often getting stuck in local minima for very long periods of time. The uniform proposal has the extreme opposite feature, as it picks a possible step at random, it has a very low acceptance rate meaning slow exploration of the landscape as the energy of a proposed state is unlikely to be close or lower. When a step is accepted however, the hamming distance is very large. Both of the above points are made clear in Figure 6.
Any CG method that does not multiple sample will have the locality issue of occasionally getting stuck in minima. If one does multiple sample, it becomes clear that the Hamming distance is slightly reduced compared to the uniform or QeMCMC method. The energy difference of proposals still outperforms classical methods, almost matching the QeMCMC. One interesting artefact is that the improved local group (without multiple sampling) greatly reduces the energy difference compared to all methods, while increasing the Hamming distance compared to the classical local method. This means that if one were to choose a problem where the local method performs well (for example in particular temperature regimes) then the improved local method is always likely to sustainably outperform the classical proposals. Scaling of this artifact is left to future work.
@@ B.2 Temperature and convergence for large systems
The relationship between temperature, system size and thermalization is conveyed through analysis of the energy convergence in Figure 7. One can see that the thermalization of the uniform proposal becomes useless for large systems, the local proposal struggles at low temperatures and in every case, both quantum proposals analysed efficiently find a better than classical low energy regime. magnetization plots are not provided as they falsely suggest convergence of non-converged distributions.
Figure 7: Average energy thermalization plot of $20$ Markov Chains at $100$ step intervals for one randomly selected Ising model of size $16$ (left), $25$ (middle) and $36$ (right). $T=0.1$ (top) is clearly a more difficult problem than $T=1$ (bottom) where the Markov chains acceptance rate is much higher. The colored dotted lines represent the $10$ lowest energy levels of the system, while the black line represents the exact average Boltzmann energy.
@@ B.3 Temperature and scaling
The temperature dependence of spectral gap scaling is briefly conveyed here through a cross-section of temperature regimes. Figure 8 shows the scaling behavior of each proposal method for $T=\{0.1,1\}$ while Table 1 shows results of each fit according to Equation 16.